OpenStudy (itsmichelle29):

What is the common difference for the sequence shown below? coordinate plane showing the points 3, 4; 1, 3; and negative 1, 2 −2 −one half one half 2

OpenStudy (solomonzelman):

For coordinates (x,y) it is equivalent of \(a_x=y\)

OpenStudy (itsmichelle29):

what does common differnce mean ??

OpenStudy (solomonzelman):

Common difference, in arithmetic sequence, is the number that you add (or subtract) to get to the next term

OpenStudy (solomonzelman):

this number is denoted with letter d.

OpenStudy (itsmichelle29):

okay thanks

OpenStudy (solomonzelman):

wow, you are very smart! genius! \(a_3=4\) \(a_1=3\) \(a_n=a_1+d(n-1)\) \(a_3=a_1+d(3-1)\) \(a_3=a_1+2d\) \(4=3+2d\) \(1=2d\) \(d=1/2\)

OpenStudy (itsmichelle29):

Lol u make me feel so much better about my math .. Thanks

OpenStudy (solomonzelman):

You should feel better about your math. I rarely meet people who catch things so quickly.

OpenStudy (solomonzelman):

you are always welcome!